FURTHER STUDIES ON SHARP TRANSITIONS INDUCED IN SOLID PARAHYDROGEN BY METHANE

\underline{C. MICHAEL LINDSAY} and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago IL, 60637, USA; TAKAMASA MOMOSE, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan.

Last year at this meeting we reported our serendipitous observation of ∼ 80 sharp transitions observed in the Q_{1}(1) region of hydrogen apparently induced by methane imbedded in the parahydrogen crystal. We have since recorded the FTIR and high-resolution laser spectra of several more methane doped parahydrogen crystals with varying methane and orthohydrogen concentrations. We are in the initial stages of analyzing this data, and it appears that there are a number of interesting phenomena occurring. In this talk we will present the new data and discuss our progress in the understanding of this fascinating problem.